OpenStudy (anonymous):

How to integrate u^2/(sqrt(a^2-u^2)

OpenStudy (anonymous):

is that supposed to be an a^2??

OpenStudy (anonymous):

Yes, it is very similar to 1\[\frac{ 1 }{ a ^{2}+x ^{2} } = \frac{ 1 }{ a }\tan^{-1} \frac{ x }{ a }\] but is has the x^2 on top

OpenStudy (john_es):

It should be interesting this, \[\int\frac{u^2}{\sqrt{a^2-u^2}}du=\int\frac{a^2}{\sqrt{a^2-u^2}}du-\int\frac{a^2-u^2}{\sqrt{a^2-u^2}}du\] And I think it should be easy from this point.

OpenStudy (anonymous):

How did you get you get that John_ES ?

OpenStudy (turingtest):

are you doing trig substitution integrals yet?

OpenStudy (anonymous):

yes

OpenStudy (turingtest):

then they probably want you to use the substitution\[u=\frac1a\sec\theta\]

OpenStudy (turingtest):

for the record, John ES used the fact that u^2=a^2-(a^2-u^2), but I don't think that is the way to go with this problem

OpenStudy (anonymous):

Thx, will try that this afternoon